# Coverage Biases in High-Resolution Satellite Imagery

### v1

@@ Conclusion
In conclusion, we find that the availability of (very) high-resolution optical satellite imagery is not equally distributed across earth. While physical phenomena such as orbital trajectories can explain parts of that bias, differences are mainly due to business considerations based on the geo-political and socio-economic situation on the ground. While levels of the digital dividend a region can reap from leveraging EO data for local development is unknown, we expect that unmet latent demand leave poorer regions behind their economic growth potential. Even though this study offers an initial glimpse into these relationships, it is not without shortcomings: First of all, we cover the major, but not all providers of (very) high-resolution satellite imagery. Although we do not expect this to change the conclusions we can draw about the existence of coverage biases, we cannot rule it out. Second, as quantifying the economic value of using satellite imagery, what we call the digital dividend of EO data in this study, is beyond the scope of this paper, we cannot assess the significance of these biases for local development. Third and related to that, since we concentrate on (very) high-resolution imagery, we leave out the most prominent and used satellite programs, namely NASA’s Landsat program and ESA’s Sentinel program, out of the study. As government-led programs, they follow a ”gotta catch ’em all” strategy, which significantly differs from the tasking-driven strategies of most privately-owned operators (with an exception of Planet). While many EO applications such as land use management, urban planning and disaster risk assessment, build on mid-resolution Landsat or Sentinel data and since these programs seem to be less affected by business model-driven coverage biases, the effect of these biases on the digital dividend might thus be less pronounced. Fourth, it is likely that our regression approach suffers from mis-specification such as omitted variable bias due to the scarcity of globally harmonized subnational data of relevance.
@@
Dependent variable: Planet image counts
@@
Dependent variable: Total image counts

### v2

@@ Conclusion
In conclusion, we find that the availability of (very) high-resolution optical satellite imagery is not equally distributed across earth. While physical phenomena such as orbital trajectories can explain parts of that bias, differences are mainly due to business considerations based on the geo-political and socio-economic situation on the ground. While levels of the digital dividend a region can reap from leveraging EO data for local development is unknown, we expect that unmet latent demand leave poorer regions – as measured by the SHDI in this study – behind their economic growth potential. Even though this study offers an initial glimpse into these relationships, it is not without shortcomings: First of all, we cover the major, but not all providers of (very) high-resolution satellite imagery. Although we do not expect this to change the conclusions we can draw about the existence of coverage biases, we cannot rule it out. Second, as quantifying the economic value of using satellite imagery, what we call the digital dividend of EO data in this study, is beyond the scope of this paper, we cannot assess the significance of these biases for local development. Third and related to that, since we concentrate on (very) high-resolution imagery, we leave out the most prominent and used satellite programs, namely NASA’s Landsat program and ESA’s Sentinel program, out of the study. As government-led programs, they follow a ”gotta catch ’em all” strategy, which significantly differs from the tasking-driven strategies of most privately-owned operators (with an exception of Planet). While many EO applications such as land use management, urban planning and disaster risk assessment, build on mid-resolution Landsat or Sentinel data and since these programs seem to be less affected by business model-driven coverage biases, the effect of these biases on the digital dividend might thus be less pronounced. Fourth, it is likely that our regression approach suffers from mis-specification such as omitted variable bias due to the scarcity of globally harmonized subnational data of relevance.
@@
Dependent variable: Historic image count (Planet only)
@@
Dependent variable: Historic image count
@@
Dependent variable: Historic image count
(1) (2) (3) (4)
Constant 0.038∗∗ -0.173∗∗∗ -0.053∗∗∗ -0.058∗∗∗
(0.016) (0.019) (0.008) (0.008)
Subnational HDI -0.034 -0.036 0.013 0.011
(0.021) (0.019) (0.008) (0.008)
Latitude (abs.) 0.100∗∗∗ 0.037∗∗∗ 0.040∗∗∗
(0.019) (0.007) (0.007)
Longitude (abs.) 0.428∗∗∗ 0.101∗∗∗ 0.101∗∗∗
(0.027) (0.011) (0.011)
# of households 0.076∗∗∗ 0.078∗∗∗
(0.006) (0.006)
Area size 0.928∗∗∗ 0.931∗∗∗
(0.010) (0.010)
Cloud coverage 0.018∗∗∗
Observations 1726 1726 1726 1726
$R^{2}$ 0.542 0.620 0.939 0.939
Adjusted $R^{2}$ 0.489 0.576 0.931 0.932
Residual Std. Error 0.036 (df=1546) 0.033 (df=1544) 0.013 (df=1542) 0.013 (df=1541)
F Statistic 10.234∗∗∗ (df=179; 1546) 13.928∗∗∗ (df=181; 1544) 129.068∗∗∗ (df=183; 1542) 129.975∗∗∗ (df=184; 1541)
Note: Country-level fixed effects included. p $<$ 0.1; ∗∗p $<$ 0.05; ∗∗∗p $<$ 0.01
Table 10: OLS with country-level fixed effects: Regressing the number of publicly available historic satellite images (without planet images) from the years 2017–23 on socio-economic and physical factors. All variables are normalized.
Dependent variable: Historic image count
(1) (2) (3) (4)
Constant -$0∗$011∗∗ -0.009∗∗∗ -0.013∗∗∗
(0.005) (0.005) (0.002) (0.002)
Income index 0.042∗∗∗ -0.006 0.009∗∗∗ 0.007∗∗
(0.007) (0.009) (0.003) (0.003)
Latitude (abs.) 0.047∗∗∗ 0.006∗∗∗ 0.009∗∗∗
(0.007) (0.002) (0.002)
Longitude (abs.) 0.054∗∗∗ 0.014∗∗∗ 0.013∗∗∗
(0.005) (0.002) (0.002)
# of households 0.075∗∗∗ 0.076∗∗∗
(0.005) (0.005)
Area size 1.034∗∗∗ 1.035∗∗∗
(0.008) (0.008)
Cloud coverage 0.008∗∗∗
Observations 1726 1726 1726 1726
$R^{2}$ 0.023 0.085 0.913 0.914
Adjusted $R^{2}$ 0.022 0.083 0.913 0.914
F Statistic 39.689∗∗∗ (df=1; 1724) 53.104∗∗∗ (df=3; 1722) 3626.481∗∗∗ (df=5; 1720) 3054.818∗∗∗ (df=6; 1719)
Note: p $<$ 0.1; ∗∗p $<$ 0.05; ∗∗∗p $<$ 0.01
Dependent variable: Historic image count ($0-0.5$ meter resolution)
(1) (2) (3) (4)
Constant -0.015∗∗ -0.007 -0.012∗∗∗ -0.013∗∗∗
(0.006) (0.006) (0.002) (0.002)
Subnational HDI 0.055∗∗∗ -0.012 0.018∗∗∗ 0.016∗∗∗
(0.008) (0.012) (0.004) (0.004)
Latitude (abs.) 0.056∗∗∗ 0.004 0.006
(0.008) (0.003) (0.003)
Longitude (abs.) 0.063∗∗∗ 0.014∗∗∗ 0.014∗∗∗
(0.006) (0.002) (0.002)
# of households 0.103∗∗∗ 0.104∗∗∗
(0.007) (0.007)
Area size 1.149∗∗∗ 1.149∗∗∗
(0.011) (0.011)
Cloud coverage 0.004
Observations 1726 1726 1726 1726
$R^{2}$ 0.024 0.085 0.885 0.885
Adjusted $R^{2}$ 0.023 0.083 0.885 0.885
Residual Std. Error 0.057 (df=1724) 0.055 (df=1722) 0.020 (df=1720) 0.020 (df=1719)
F Statistic 42.491∗∗∗ (df=1; 1724) 53.273∗∗∗ (df=3; 1722) 2648.491∗∗∗ (df=5; 1720) 2209.375∗∗∗ (df=6; 1719)
Note: p $<$ 0.1; ∗∗p $<$ 0.05; ∗∗∗p $<$ 0.01
Table 12: OLS: Regressing the number of publicly available historic satellite images with a resolution $0-0.5$m from the years 2017–23 on socio-economic and physical factors. All variables are normalized.
@@
Figure 14: Lorenz curve showing the distribution of historic satellite image availability (images per km²) across regions, ordered by their Subnational Human Development Index (SHDI). The curve deviates notably from the dotted line of equality, indicating a higher concentration of high-resolution imagery per unit area in more developed regions. The Gini coefficient is 0.638.